OpenStudy (anonymous):

What is the eccentricity of the ellipse shown below? (x-5)^2/9+(y+2)^2/25=1

OpenStudy (anonymous):

\[\frac{ \sqrt{b^2-a^2} }{b }\]

OpenStudy (anonymous):

b^2=25 a^2=9

OpenStudy (anonymous):

tese are my options

OpenStudy (anonymous):

I think you're on the wrong problem

OpenStudy (anonymous):

oh! i mistyped the problem!

OpenStudy (anonymous):

its (x-5)^2/52+(y+1)^2/64=1

OpenStudy (anonymous):

same idea.. plug the values into that expression I gave...

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

yes